# Two-phase formation of galaxies: the coevolution between galaxies and dark matter halos

### v2

@@ 4.2 The origin of dynamical hotness: concentrated versus scattered star formation
As suggested by numerical simulations (e.g. Keres et al., 2005; Dekel & Birnboim, 2006; Dekel et al., 2009; Danovich et al., 2015), galaxies at high $z$ are often fed by cold streams of gas that are seeded by large-scale dark-matter filaments and condensed by radiative cooling. Scattered star formation is thus expected to take place in these cold streams. This is indeed the case for FIRE-2 galaxies, as demonstrated by the example shown in Fig. 7(c). The environment provided by a fast-accreting halo at high $z$ can thus be understood as to provide a condition for the IGM/CGM similar to that satisfied in satellite galaxies in massive halos at low $z$, where gas density is sufficiently high to allow local self-gravitation and star formation. Note that the large extents and filamentary structures of newly formed stars also predict features in the size and morphology of high-$z$ galaxies directly testable by deep observations, as we will discuss in §5.
@@
Given the two distinct modes of star formation that build up the dynamically hot stellar component, a question arises as what drives the transition of galaxies from the scattered star formation to the concentrated star formation. This again can be inferred from the evolution of galaxies in the quadrant diagram. An obvious difference between the two regimes is the assembly rates of the host halos. As suggested by N-body simulations (Borzyszkowski et al., 2017), fast-accreting halos are often connected with filaments in the cosmic web, while slow-accreting halos are not. The cold streams feeding the formation of bulges in the earlier regime ($\gamma\gtrsim 3/8$) are thus expected to disappear in the later regime ($\gamma\approx 0$$3/8$), as dark matter filaments seeding the gaseous streams disappear. Consequently, the high density condition for scattered star formation in CGM/IGM vanishes in the later regime, allowing the gas to accumulate into the halo center without being consumed by star formation and feedback until the condition of self-gravitation is satisfied globally to trigger an intensive and concentrated star formation. During the transitional phase, the potential well of the host halo, as quantified by the maximum circular velocity, $V_{\rm max}$, has already reached the peak value among all redshifts (Zhao et al., 2003). This, together with the rich amount of gas accumulated during the fast phase and the self-gravity of the gas that further deepens the potential well, creates a condition that is highly resistant to feedback. A large amount of gas can thus be converted into stars in a short period, which explains the rapid growth of bulge mass during the concentrated star formation.
@@
Another mechanism that may also contribute to the transition from the scattered to the concentrated star formation is ‘dynamical heating’, by which stars formed at early epochs are ‘heated’ later on by gravitational perturbations. One source of dynamical heating is the fast assembly of the dark matter halo, whose effect is expected to be strong at high $z$. The large extents of bulges formed via scattered star formation at high $z$ imply that the total gravity on the bulge is dominated by that of the dark matter. As halos in the fast phase self-similarly expand its structure (Wechsler et al., 2002; Zhao et al., 2003), stars co-expand with dark matter as they gain energy from the fast-varying gravitational potential. In Fig. 6(a), we show the evolution of the half-mass radius of the stars formed at $z\geqslant 8$ (dark-blue curve). The half-mass radius of these old stars expands with time, suggesting that dynamical heating is indeed effective. The long-term trend of such an expansion follows that of the halo radius, indicating that dynamical heating by fast halo assembly is likely an important heating mechanism. The other sources of dynamical heating can be from episodic outflow produced by bursty star formation that causes temporary loss of binding energy of stars, and the formation of dense gas/stellar clumps that scatter stars and transfer energy to them. Such effects have been suggested by e.g. El-Badry et al. (2016), who showed that bursty star formation leads to outward migrations of stars, more so for older stars (Kado-Fong et al., 2022), and by McCluskey et al. (2024), who demonstrated that the velocity dispersion of stars rapidly increases with time at high $z$. Separating the sources of dynamical heating, however, is not straightforward, as both dark-matter and baryonic sources of dynamical heating are expected to be more effective at higher $z$. Future studies utilizing idealized simulations can help to disentangle these sources, and to quantify their relative importance.
@@
While the transition between the two regimes of bulge formation is clearly seen in our sample of galaxies, we note that this largely relies on the specific sub-grid models adopted in the FIRE-2 simulations: once gas is accreted into a halo through filaments and condensed to a density sufficient for star formation, the ‘early’ feedback produced by young stars through, e.g. stellar winds and ionizing radiation, may be able to disperse the gas before more stars can form. Such feedback from star formation has been investigated in some hydrodynamic simulations for isolated galaxies (Übler et al., 2014; Agertz & Kravtsov, 2015, 2016; Li et al., 2020; Deng et al., 2024), and is expected to be important also in cold filaments due to their shallower gravitational potential well than that in the ISM. In FIRE-2, the early feedback was included in the form of internal radiative feedback from the stars and external radiative feedback from the UV background, but the implemented feedback strength may not be sufficient to prevent star formation in the filaments. Future observations able to detect the low-surface-brightness outskirts of high-$z$ galaxies are crucial to test the presence of the scattered star formation and to narrow down sub-grid models to be implemented in simulations.
@@ 4.3 Processes driving the transition to cold disk
As discussed above in connection to the quadrant diagram, the formation of a thin disk requires two conditions. First, the gravitational potential well should be stable so that dynamical heating by variation of the gravitational potential is not important. A high accretion rate of the halo, represented by a high value of $\gamma$, is a primary driver of dynamical heating on galactic scales. Thus, the formation of thin disks is expected only when $\gamma$ is low, as is demonstrated clearly by the strong correlation between $t_{\rm thin}$ and $t_{\gamma}$ in Fig. 3. The second condition is that the cold gas fraction must be sufficiently low so that the total gravitational potential in the inner region of the halo is not dominated by the disk. This is demonstrated by the correlation between $f_{\rm thin}$ and $f_{\lambda}$ shown in Fig. 4(g), and by the rapid decline of $f_{\text{gas+star}}$ within inner $1\,{\rm{kpc}}$ at $t>t_{\gamma}^{({\rm end})}$, shown in Fig. 5(b), when the star formation starts to be dominated by that in the thin disk (see Fig. 2).
@@
In the simulated galaxies, the gravitationally dominant component undergoes two transitions. Initially, at $t<t_{\gamma}^{({\rm start})}$ when $\gamma$ is large, dark matter dominates the gravitational potential almost everywhere in the halo, as implied by the large sizes of galaxies with respect to the sizes of halos (see Fig. 6c). During this phase, a small amount of stars is formed in the ‘scattered mode’ discussed in §4.2, and the gravitational potential near the halo center is still dominated by dark matter. However, as the evolution moves to the transitional phase of halo assembly, $t_{\gamma}^{({\rm start})}<t<t_{\gamma}^{({\rm end})}$, a large amount of low-spin gas accumulates in the halo center, becomes self-gravitating, forms bulge stars in the ‘concentrated mode’, and reduce the gas by feedback. As shown in Fig. 5(b), in almost all simulated galaxies, the gravitational potential in the central part of the halo becomes dominated by star-forming gas and young stars in the bulge in this phase. As soon as the bulge component is established at $t>t_{\rm bulge,1/2}\sim t_{\gamma}$, it provides another source to stabilize the gravitational potential and to promote the formation of a dynamically cold disk from gas with sufficiently high spin (see also Hopkins et al., 2023). As the disk grows, the bulge component can continue to grow slowly from low-spin gas near the center, as shown in Fig. 1f. Thus, the transition to thin-disk formation in the simulated galaxies is well described by the moving of a galaxy to Q3 in the quadrant diagram. In particular, the formation of a dynamically hot bulge in the fast accretion regime of the halo promotes the formation of a thin disk by stabilizing the gravitational potential as well as by reducing the amount of low-spin gas, making the transition to thin-disk formation at $t>t_{\gamma}^{({\rm end})}$ more prominent.
@@
Figure 9: Distribution of stellar bulges in the $b/a$-$c/a$ plane. Here we only include young bulge stars in given ranges of galactocentric distance: (a), $r<0.25R_{\rm 90,*}$; (b), $0.25R_{\rm 90,*}<r<0.5R_{\rm 90,*}$. Data points show the snapshots of all galaxies in our sample, with color coded according to the redshift smoothed by the LOESS method (Cleveland & Devlin, 1988). In each panel, squares show the evolution of the axis ratios obtained by the median of all galaxies. This figure verifies that scattered star formation in high-$z$ galaxies takes place at large radii in cold streams, thus producing anisotropic shapes of bulges. See §5 for details.
@@ 5 Structural evolution of dynamically hot stellar systems
A particularly interesting part of the two-phase formation of galaxies is the growth of dynamically hot stellar systems (referred to as bulges in this paper) in fast-accreting halos. The importance of bulge formation is threefold. First, the ‘scattered formation’ of stars in the early epoch depends on the competition between various cooling and heating processes. Many of these processes are unresolved in current simulations and are implemented as subgrid models. The bulge components formed through this channel may thus encode information about these processes, and their properties may be used to distinguish models. Second, as shown in Fig. 4(e) and other previous studies using different simulations (Pillepich et al., 2019; Yu et al., 2023), the bulge components dominate the total stellar mass of high-$z$ galaxies, and thus their properties have important implications for observations targeting high-$z$ galaxies. Recent observations with resolved stellar dynamics reveal surprising cases of disk-like galaxies at $z\gtrsim 4$, which appears at odd with theoretical expectations. Third, bulges are found to co-evolve with the central supermassive black holes (SMBHs) of galaxies (Kormendy & Ho, 2013; Greene et al., 2020; Zhuang & Ho, 2023), and thus closely related to understanding galaxy quenching (Dekel & Burkert, 2014; Woo et al., 2015; Lin et al., 2019; Hong et al., 2023; Mo et al., 2024). The bulges in FIRE-2 may thus be used to study potential sites of SMBH growth. Here we focus on bulges in FIRE-2 by identifying properties that not only reflect the physical processes driving the formation and evolution, but also can be measured observationally.
Within the two-phase model, Chen et al. (2025b) showed that the fraction of star formation in dense star clusters is expected to increase towards high $z$ (Chen et al., 2025b, see their fig. 18), and such a trend can be used to explain the formation of old globular clusters. To see if this expectation is supported by FIRE-2 galaxies, we show, in Fig. 7(a)–(c), the projected stellar distribution of the bulge component of the galaxy m12f at three redshifts chosen to represent different stages of bulge formation. The change in the clumpiness with time is obvious: the galaxy is highly clumpy at $z=4.4$, becomes less clumpy at $z=1.5$, and appears smooth at $z=0$. A similar finding was also reported by Ma et al. (2020) for a number of other galaxies in the FIRE-2 suite.
@@
The evolution of the bulge structure can be seen from the example images shown in Fig. 7(a)–(c). At $z=4.4$, the stellar distribution is elongated along the diagonal direction of the image, exhibiting filament-like structures that trace the accretion flow of gas into the galaxy. We visually inspected the cold gas stream and found that it is indeed aligned with the stellar distribution. Meanwhile, the stellar distribution is extended. The size of the bulge (indicated by a dashed circle; see also Fig. 6a) is already comparable to the size at $z=0$, although the stellar mass is much lower than the final bulge mass (see Fig. 2a). The concentration of stars in the galaxy is evident, but the multiple dense stellar clumps near the center make it difficult to identify the center. At $z=1.5$, the stellar distribution is less elongated, with a well-defined concentration. At $z=0$, the bulge looks spherical, and the concentrated core is more prominent. In panels (d)–(f), we show the evolution of the stellar mass surface density profile $\Sigma_{*}(r_{\rm p})$], the Sérsic index ($n$), and the shape parameters (axis ratio, $c/a$, and the triaxiality, $T$), respectively, of this bulge. As can be seen, the bulge evolves from a prolate, low-concentration structure at high $z$ to a spheroidal, high-concentration structure at low $z$. There appears to be a clear transition in the evolution of the structural parameters, and the transition time corresponds well to that in halo assembly ($t_{\gamma}$, marked by the red solid lines in panels e and f) and in dynamical hotness of the galaxy ($t_{\cal H}$, marked by the red dashed lines in panels e and f). This provides further support that the evolution of bulges is driven by the evolution of dark matter halos.
@@
In Fig. 8, we show the evolution of the bulge structure for all the 14 galaxies in our example. In panels (a)–(d), we show the evolution of the structural parameters: Sérsic index $n$, triaxiality $T$, $b/a$ and $c/a$, respectively, as functions of cosmic time for the median and $25^{\rm th}$$75^{\rm th}$ percentiles among all bulges. All these results confirm the expectation that the bulge structure evolves from being prolate (small $c/a$ and $b/a$; large $T$) to being spheroidal (large $c/a$ and $b/a$; small $T$), and the bulge concentration (Sérsic index $n$) increases with time. The scatter of $n$ is large at high $z$, reflecting the presence of violent dynamical disturbances that can distort the distribution of stars and prevent the formation of a well-defined galactic center (see also Fig. 7c and f). We checked the Sérsic fitting and found that the $\chi^{2}$ is significantly enlarged at $z\gtrsim 2$, again indicating the distorted stellar distribution and the difficulty of using a single analytical profile to capture such distortion. The cosmic times in panels (a)–(d) are aligned with $t_{\gamma}^{\rm(start)}$ for individual galaxies. Clearly, the transition in shape properties is closely related to the transition of the halo assembly. This tight relation between the transition in bulge shape and the transition in halo assembly seems to indicate that the accretion rate and accretion configuration are correlated with each other, so that the change from fast to slow accretion of the halo corresponds to the change from an anisotropic cold-mode accretion to an isotropic hot-mode accretion of gas into the galaxy. In panel (e), we show the locations of bulges in the c/ac/ab/ab/a plane at two epochs: one is the earliest time when a reliable measurement of the shape parameters can be made for a galaxy (when the number of the star particles is larger than 1000) in the dynamically hot phase, and the other is at $z=0$, when the halo is already in slow accretion. We see clearly that the evolution is from being prolate to being spheroidal (see also Meng & Gnedin, 2021; Xia et al., 2025).
@@
In observations, resolving structural properties such as the shape and concentration is less demanding than resolving small-scale clumpiness. The evolution of the bulge structure thus provides a promising way to test the two-phase formation model by observations. Indeed, recent observations of high-$z$ galaxies with JWST indicate an excessively large fraction of galaxies with prolate stellar distribution (Law et al., 2012; Yuma et al., 2012; van der Wel et al., 2014b; Zhang et al., 2019; Pandya et al., 2024). However, the small Sérsic index shown in Fig. 8(a) for FIRE-2 galaxies at high $z$ indicates that the measurements of the Sérsic index may not be able to distinguish prolate ellipsoids from inclined disks. Caution is thus needed when interpreting the density profiles of high-$z$ galaxies. For comparison, we show the evolution of the median of individual structural parameters using all stars in Fig. 8(a)–(d). Except for the Sérsic index, the evolution of the structural parameters is less significant than that obtained using only bulge stars, suggesting that a precise decomposition between the bulge and disk components is needed to constrain the formation mechanism of dynamically hot stellar systems.

### v3

@@ 4.2 The origin of dynamical hotness: concentrated versus scattered star formation
As suggested by numerical simulations (e.g. Keres et al., 2005; Dekel and Birnboim, 2006; Dekel et al., 2009; Danovich et al., 2015), galaxies at high $z$ are often fed by cold streams of gas that are seeded by large-scale dark-matter filaments and condensed by radiative cooling. Scattered star formation is thus expected to take place in these cold streams. This is indeed the case for FIRE-2 galaxies, as demonstrated by the example shown in Fig. 8(c). The environment provided by a fast-accreting halo at high $z$ can thus be understood as to provide a condition for the IGM/CGM similar to that satisfied in satellite galaxies in massive halos at low $z$, where gas density may be sufficiently high to allow local self-gravitation and star formation. Note that the large extents and filamentary structures of newly formed stars also predict features in the size and morphology of high-$z$ galaxies directly testable by deep observations, as we will discuss in §5.
@@
Given the two distinct modes of star formation that build up the dynamically hot stellar component, a question arises as what drives the transition of galaxies from the scattered star formation to the concentrated star formation. This again can be inferred from the evolution of galaxies in the quadrant diagram. An obvious difference between the two regimes is the assembly rate of the host halo. As suggested by N-body simulations (Borzyszkowski et al., 2017), fast-accreting halos are often connected with filaments in the cosmic web, while slow-accreting halos are not. The cold streams feeding the formation of bulges in the earlier regime ($\gamma\gtrsim 3/8$) are thus expected to disappear in the later regime ($\gamma\approx 0$$3/8$), as dark matter filaments seeding the gaseous streams disappear. Consequently, the high density condition for scattered star formation in CGM/IGM may not be satisfied in the later regime, allowing the gas to accumulate into the halo center without being consumed by star formation and feedback until the condition of self-gravitation is satisfied globally to trigger an intense and concentrated star formation. During the transitional phase, the potential well of the host halo, as quantified by the maximum circular velocity, $V_{\rm max}$, has already reached the peak value among all redshifts (Zhao et al., 2003). This, together with the rich amount of gas accumulated during the fast phase and the self-gravity of the gas that further deepens the potential well, creates a condition that is highly resistant to feedback (Hopkins et al., 2023). A large amount of gas can thus be converted into stars in a short period, which explains the rapid growth of bulge mass during the concentrated star formation.
@@
Another mechanism that may also contribute to the transition from the scattered to the concentrated star formation is ‘dynamical heating’, by which stars formed at early epochs are ‘heated’ later on by gravitational perturbations. One source of dynamical heating is the fast assembly of the dark matter halo, whose effect is expected to be strong at high $z$. The large extents of bulges formed via scattered star formation at high $z$ imply that the total gravity on the bulge is dominated by that of the dark matter. As halos in the fast phase self-similarly expand their structure (Wechsler et al., 2002; Zhao et al., 2003), stars co-expand with dark matter as they gain energy from the fast-varying gravitational potential. In Fig. 7(a), we show the evolution of the half-mass radius of the stars formed at $z\geqslant 8$ (dark-blue curve). The half-mass radius of these old stars expands with time, suggesting that dynamical heating is indeed effective. The long-term trend of such an expansion follows that of the halo radius, indicating that dynamical heating by fast halo assembly is likely an important heating mechanism. The other sources of dynamical heating can be from episodic outflow produced by bursty star formation that causes temporary loss of binding energy of stars, and the formation of dense gas/stellar clumps that scatter stars and transfer energy to them. Such effects have been suggested by e.g. El-Badry et al. (2016), who showed that bursty star formation leads to outward migrations of stars, more so for older stars (Kado-Fong et al., 2022), and by McCluskey et al. (2024), who demonstrated that the velocity dispersion of stars rapidly increases with time at high $z$. Separating the sources of dynamical heating, however, is not straightforward, as both dark-matter and baryonic sources of dynamical heating are expected to be more effective at higher $z$. Future studies utilizing idealized simulations can help to disentangle these sources, and to quantify their relative importance.
@@
While the transition between the two regimes of bulge formation is clearly seen in our sample of galaxies, we note that this largely relies on the specific sub-grid models adopted in the FIRE-2 simulations: once gas is accreted into a halo through filaments and condensed to a density sufficient for star formation, the ‘early’ feedback produced by young stars through, e.g. stellar winds and ionizing radiation, may be able to disperse the gas before more stars can form. Such feedback from star formation has been investigated in some hydrodynamic simulations for isolated galaxies (Übler et al., 2014; Agertz and Kravtsov, 2015, 2016; Li et al., 2020; Deng et al., 2024), and is expected to be important also in cold filaments due to their shallower gravitational potential well than that in the ISM. In FIRE-2, the early feedback was included in the form of internal radiative feedback from the stars and external radiative feedback from the UV background, but the implemented feedback strength may not be sufficient to prevent star formation in the filaments. Future observations able to detect the low-surface-brightness outskirts of high-$z$ galaxies are crucial to test the presence of the scattered star formation and to narrow down sub-grid models to be implemented in simulations.
While fast halo accretion is related to bulge formation in our sample of FIRE-2 galaxies, it should not be considered as a necessary condition for galaxies to be dynamically hot. Other “heating” sources may also disturb galaxies and contribute to the dynamical hotness. For example, galaxy-galaxy mergers and close encounters were found to be critical in shaping the structure of massive elliptical galaxies at late times (e.g. Shen et al., 2003; Naab et al., 2009; Kluge and Bender, 2023); the changes in the gravitational potential caused by inflow/outflow in intense star-formation cycles were proposed as a mechanism to produce the structural diversity of dwarf galaxies (e.g. El-Badry et al., 2016; Di Cintio et al., 2017); rapid and/or filamentary gas accretion can provide additional dynamical heating beside the gravity of dark matter (Dekel and Birnboim, 2006; Robertson and Goldreich, 2012; Ginzburg et al., 2022; Goldner et al., 2025); more exotic heating mechanisms, such as dark-matter self-interaction (Balberg et al., 2002; Zhang et al., 2025), may also produce dynamical hotness. The fast halo accretion is neither a sufficient condition for galaxies to be dynamically hot. As suggested by Mo et al. (2024), one additional condition for the formation of dynamically hot galaxies is that gas cooling in the CGM is rapid so as to sustain gravitational instability and active star formation. In Appendix B, we show that the cooling time is indeed shorter than the free-fall time during the fast-accretion phase for galaxies in our sample, so that the condition of rapid cooling is satisfied.
@@ 4.3 Processes driving the transition to cold disk
As discussed above in connection to the quadrant diagram, the formation of a thin disk requires two conditions. First, the gravitational potential well should be stable so that dynamical heating by variation of the gravitational potential is not important. A high accretion rate of the halo, represented by a high value of $\gamma$, can perturb the gravitational potential and dynamically heat baryons on galactic scales. Thus, the formation of thin disks occurs predominantly when $\gamma$ is low, consistent with the correlation between $t_{\rm thin}$ and $t_{\gamma}$ shown in Fig. 3. The second condition is that the cold gas fraction must be sufficiently low so that the total gravitational potential in the inner region of the halo is not dominated by the disk. This is demonstrated by the correlation between $f_{\rm thin}$ and $f_{\lambda}$ shown in Fig. 5(g), and by the rapid decline of $f_{\text{gas+star}}$ within inner $1\,{\rm{kpc}}$ at $t>t_{\gamma}^{({\rm end})}$, shown in Fig. 6(b), when the star formation starts to be dominated by that in the thin disk (see Fig. 2).
@@
In the simulated galaxies, the gravitationally dominant component undergoes two transitions. Initially, at $t<t_{\gamma}^{({\rm start})}$ when $\gamma$ is large, dark matter dominates the gravitational potential almost everywhere in the halo, as implied by the large sizes of galaxies with respect to the sizes of halos (see Fig. 7c). During this phase, a small amount of stars is formed in the ‘scattered mode’ discussed in §4.2, and the gravitational potential near the halo center is still dominated by dark matter. However, as the evolution moves to the transitional phase of halo assembly, $t_{\gamma}^{({\rm start})}<t<t_{\gamma}^{({\rm end})}$, a large amount of low-spin gas accumulates in the halo center, becomes self-gravitating, forms bulge stars in the ‘concentrated mode’, and reduces the gas by feedback. As shown in Fig. 6(b), in almost all simulated galaxies, the gravitational potential in the central part of the halo becomes dominated by star-forming gas and young stars in the bulge in this phase. As soon as the bulge component is established at $t>t_{\rm bulge,1/2}\sim t_{\gamma}$, it provides another source to stabilize the gravitational potential and to promote the formation of a dynamically cold disk from gas with sufficiently high spin (see also Hopkins et al., 2023). As the disk grows, the bulge component can continue to grow slowly from low-spin gas near the center, as shown in Fig. 1f. Thus, the transition to thin-disk formation in the simulated galaxies is well described by the moving of a galaxy to Q3 in the quadrant diagram. In particular, the formation of a dynamically hot bulge in the fast accretion regime of the halo promotes the formation of a thin disk by stabilizing the gravitational potential as well as by reducing the amount of low-spin gas, making the transition to thin-disk formation at $t>t_{\gamma}^{({\rm end})}$ more prominent.
As described in the end of §2, the three galaxies, m12w, m12z and m12r, are exceptional in their transitions in galaxy dynamics and halo assembly. These exceptions point to the presence of additional heating sources, such as those discussed at the end of §4.2, that can delay the dynamical settling and therefore deserve special attention. Galaxies m12w and m12z maintain a high gas fraction after the transitional phase of halo assembly (see the pink triangles and purple squares in Fig. 6b). A more detailed examination shows that their values of $f_{\lambda}$ are close to $1$ at $z\approx 0$. These two galaxies thus never entered Q3 of the quadrant diagram, implying that the condition of gas self-gravitating is not removed and the growth of a dynamically cold disk is prohibited. m12r is similar to these two galaxies in that it also maintains a high gas fraction after the transitional phase of halo assembly. The difference is that the halo of m12r made a transition to the slow phase very early and has already entered Q3 at $z\approx 0$. The disk of m12r is therefore capable of growing and dominating the stellar mass at $z=0$. Interestingly, the host halo of m12r has a recent major merger (with a halo mass ratio of about $1:3$) at $z\approx 0.3$, which brings into the system a satellite galaxy that merges with m12r at $z\approx 0.22$. This satellite has an orbit almost co-rotating with the disk of m12r, which allows the disk of m12r to survive the merger. At $z\approx 0.1$, another halo merges into the host halo of m12r (with a halo mass ratio of about $1:7$). This introduces a close companion to m12r at $z\approx 0$ and a galaxy-galaxy merger may occur in the near future. A question remains as to how these three exceptions maintain a high gas fraction after the transition of halo assembly. The distribution of these galaxies in the assembly-environment planes (Fig. 11) shows that they are among the galaxies with the lowest environmental density and that m12r has an exceptionally late $t_{\rm h,1/2}$. Furthermore, the halo masses of these galaxies are also relatively low (see Table 1). These properties may result in different gas inflow and outflow in FIRE-2, so that a high gas fraction can be sustained even at late times.
@@
Figure 10: Distribution of stellar bulges in the $b/a$-$c/a$ plane. Here we only include young bulge stars in given ranges of galactocentric distance: (a), $r<0.25R_{\rm 90,*}$; (b), $0.25R_{\rm 90,*}<r<0.5R_{\rm 90,*}$. Data points show the snapshots of all galaxies in our sample, with color coded according to the redshift smoothed by the LOESS method (Cleveland and Devlin, 1988). In each panel, squares show the evolution of the axis ratios obtained by the median of all galaxies. This figure verifies that scattered star formation in high-$z$ galaxies takes place at large radii in cold streams, thus producing anisotropic shapes of bulges. See §5 for details.
@@ 5 Structural evolution of dynamically hot stellar systems
A particularly interesting part of the two-phase formation of galaxies is the growth of dynamically hot stellar systems (referred to as bulges in this paper) in fast-accreting halos. The importance of bulge formation is threefold. First, the ‘scattered formation’ of stars in the early epoch depends on the competition between various cooling and heating processes. Many of these processes are unresolved in current simulations and are implemented as subgrid models. The bulge components formed through this channel may thus encode information about these processes, and their properties may be used to distinguish models. Second, as shown in Fig. 5(e) and other previous studies using different simulations (Pillepich et al., 2019; Yu et al., 2023), the bulge components dominate the total stellar mass of high-$z$ galaxies, and thus their properties have important implications for observations targeting high-$z$ galaxies. Recent observations with resolved stellar dynamics reveal surprising cases of disk-like galaxies at $z\gtrsim 4$, which appears at odd with theoretical expectations. Third, bulges are found to co-evolve with the central supermassive black holes (SMBHs) of galaxies (Kormendy and Ho, 2013; Greene et al., 2020; Zhuang and Ho, 2023), and thus closely related to understanding galaxy quenching (Dekel and Burkert, 2014; Woo et al., 2015; Lin et al., 2019; Hong et al., 2023; Mo et al., 2024). The bulges in FIRE-2 may thus be used to study potential sites of SMBH growth. Here we focus on bulges in FIRE-2 by identifying properties that not only reflect the physical processes driving the formation and evolution, but also can be measured observationally.
Within the two-phase model, Chen et al. (2025a) showed that the fraction of star formation in dense star clusters is expected to increase towards high $z$ (Chen et al., 2025a, see their fig. 18), and such a trend can be used to explain the formation of old globular clusters. To see if this expectation is supported by FIRE-2 galaxies, we show, in Fig. 8(a)–(c), the projected stellar distribution of the bulge component of the galaxy m12f at three redshifts chosen to represent different stages of bulge formation. The change in the clumpiness with time is obvious: the galaxy is highly clumpy at $z=4.4$, becomes less clumpy at $z=1.5$, and appears smooth at $z=0$. A similar finding was also reported by Ma et al. (2020) for a number of other galaxies in the FIRE-2 suite.
@@
The evolution of the bulge structure can be seen from the example images shown in Fig. 8(a)–(c). At $z=4.4$, the stellar distribution is elongated along the diagonal direction of the image, exhibiting filament-like structures that trace the accretion flow of gas into the galaxy. We visually inspected the cold gas stream and found that it is indeed aligned with the stellar distribution. Meanwhile, the stellar distribution is extended. The size of the bulge (indicated by a dashed circle; see also Fig. 7a) is already comparable to the size at $z=0$, although the stellar mass is much lower than the final bulge mass (see Fig. 2a). The concentration of stars in the galaxy is evident, but the multiple dense stellar clumps near the center make it difficult to identify the center. At $z=1.5$, the stellar distribution is less elongated, with a well-defined concentration. At $z=0$, the bulge looks spherical, and the concentrated core is more prominent. In panels (d)–(f), we show the evolution of the stellar mass surface density profile $\Sigma_{*}(r_{\rm p})$], the Sérsic index ($n$), and the shape parameters (axis ratio, $c/a$, and the triaxiality, $T$), respectively, of this bulge. As can be seen, the bulge evolves from a prolate, low-concentration structure at high $z$ to a spheroidal, high-concentration structure at low $z$. There appears to be a clear transition in the evolution of the structural parameters, and the transition time corresponds well to that in halo assembly ($t_{\gamma}$, marked by the red solid lines in panels e and f) and in dynamical hotness of the galaxy ($t_{\cal H}$, marked by the red dashed lines in panels e and f). This provides further support that the evolution of bulges is consistent with being influenced by the evolution of dark matter halos.
@@
In Fig. 9, we show the evolution of the bulge structure for all the 14 galaxies in our sample. In panels (a)–(d), we show the evolution of the structural parameters: Sérsic index $n$, triaxiality $T$, $b/a$ and $c/a$, respectively, as functions of cosmic time for the median and $25^{\rm th}$$75^{\rm th}$ percentiles among all bulges. All these results confirm the expectation that the bulge structure evolves from being prolate (small $c/a$ and $b/a$; large $T$) to being spheroidal (large $c/a$ and $b/a$; small $T$), and the bulge concentration (Sérsic index $n$) increases with time. The scatter of $n$ is large at high $z$, reflecting the presence of violent dynamical disturbances that can distort the distribution of stars and prevent the formation of a well-defined galactic center (see also Fig. 8c and f). We checked the Sérsic fitting and found that the $\chi^{2}$ is significantly enlarged at $z\gtrsim 2$, again indicating the distorted stellar distribution and the difficulty of using a single analytical profile to capture such distortion. The cosmic times in panels (a)–(d) are aligned with $t_{\gamma}^{\rm(start)}$ for individual galaxies. Clearly, the transition in shape properties is closely related to the transition of the halo assembly. This tight relation between the transition in bulge shape and the transition in halo assembly seems to indicate that the accretion rate and accretion configuration are correlated with each other, so that the change from fast to slow accretion of the halo corresponds to the change from an anisotropic cold-mode accretion to an isotropic hot-mode accretion of gas into the galaxy. In panel (e), we show the locations of bulges in the c/ac/ab/ab/a plane at two epochs: one is the earliest time when a reliable measurement of the shape parameters can be made for a galaxy (when the number of the star particles is larger than 1000) in the dynamically hot phase, and the other is at $z=0$, when the halo is already in slow accretion. We see clearly that the evolution is from being prolate to being spheroidal (see also Meng and Gnedin, 2021; Xia et al., 2025).
@@
In observations, resolving structural properties such as the shape and concentration is less demanding than resolving small-scale clumpiness. The evolution of the bulge structure thus provides a promising way to test the two-phase formation model by observations. Indeed, recent observations of high-$z$ galaxies with JWST indicate an excessively large fraction of galaxies with prolate stellar distribution (Law et al., 2012; Yuma et al., 2012; van der Wel et al., 2014a; Zhang et al., 2019; Pandya et al., 2024). However, the small Sérsic index shown in Fig. 9(a) for FIRE-2 galaxies at high $z$ indicates that the measurements of the Sérsic index may not be able to distinguish prolate ellipsoids from inclined disks. Caution is thus needed when interpreting the density profiles of high-$z$ galaxies. For comparison, we show the evolution of the median of individual structural parameters using all stars in Fig. 9(a)–(d). Except for the Sérsic index, the evolution of the structural parameters is less significant than that obtained using only bulge stars, suggesting that a precise decomposition between the bulge and disk components is needed to constrain the formation mechanism of dynamically hot stellar systems.